[Ncep.nhc.nco_contacts] URGENT - MAJOR ISSUES - Please Read!: Re: TSB Morning Rounds - Wed Jul 13, 2016
Monica Bozeman - NOAA Federal
monica.bozeman at noaa.gov
Wed Jul 13 16:03:13 UTC 2016
SDM says that data is starting to come back now.
National Hurricane Center
monica.bozeman at noaa.gov
On Wed, Jul 13, 2016 at 11:53 AM, Monica Bozeman - NOAA Federal <
monica.bozeman at noaa.gov> wrote:
> Doing some quick investigation of our data, we still seem to be getting in
> LIMITED observations, some is missing but we still have a few observations.
> Model data and other data sources that we get via dbnet were unaffected by
> the outage.
> NCO may have been able to switch over to the backup location, or we were
> just really lucky. NCWCP is still having issues.
> Looks like the SDM was just able to get a message out, so the availability
> of sending products seems to be improving.
> Monica Bozeman
> National Hurricane Center
> monica.bozeman at noaa.gov
> On Wed, Jul 13, 2016 at 11:41 AM, Monica Bozeman - NOAA Federal <
> monica.bozeman at noaa.gov> wrote:
>> Fresh off the presses! MAJOR impacts to the NWS at the moment!! Please
>> - WAN at Silver Spring is down for TOC and NCF - NCO and OpsNet are
>> working the problem. Occured at 10:45 am.
>> - No observations or SBN data will be coming in until this has been
>> fixed. Ex no METARS, no Radar
>> - NWS cannot send out any text products, radiofax charts have no
>> where to go since the TOC is down
>> - SDM can't send admin messages since their messages go through the
>> TOC and NCF.
>> NCO is aware of the issues and working to resolve them.
>> I'll keep you all posted if I hear anything from NCWCP for any updates to
>> the situation if no one is able to send out a global email update.
>> Monica Bozeman
>> National Hurricane Center
>> monica.bozeman at noaa.gov
>> On Wed, Jul 13, 2016 at 11:32 AM, TSB Admin <nhc.tsbadmin at noaa.gov>
>>> Here is today's summary of NHC computer operations:
>>> --- Monica
>>> -The TAFB shift log scripts are intermittently reporting issues to the
>>> TSB spreadsheet. TSB will investigate.
>>> -TAFB Dvorak fixes came in faster overnight. TSB will continue to
>>> monitor this issue in case additional tweaks are needed again.
>>> -Himawari data is still out. TSB will investigate.
>>> -TSB is still investigating duplicate shift log emails.
>>> -TSB is still investigating the Surface Wind Field graphics issues.
>>> *Other items:*
>>> -The Storm Surge AWIPS Sprint continues all week.
>>> -After the HWRF/GFDL upgrades yesterday, the grib2 files are arriving
>>> later than usual. Note that the new files are much larger than the old
>>> files, which may account for the latency. TSB is working with NCO to
>>> resolve this issue.
>>> *NHC RFCs:*
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts